What is the length of the same rail road track on the cold winter day when the temperature is 0 degree Celsius


1
Expert's answer
2021-04-22T10:59:51-0400

We know that

Length expansion formula

"L=L_{0}(1+\alpha\Delta T)"

"L=L_{0}(1+\alpha(T_{0}-T_{1}))\rightarrow(1)"

Where

"L_{0}=" Initial length

"\alpha" =linear expansion cofficient

"T_{1}=" Final temperature


"T_{0}=" initial temperature

Put T0=0°c

"L=L_{0}(1+\alpha(0-T_{1}))"

"L=L_{0}(1-\alpha T_{1})"



"(L-L_{0})=-L_{0}\alpha T_{1}"

"\frac{L-L_{0}}{L_{0}}=-\alpha T_{1}\rightarrow(2)"

equation (2) repersent rail road track length is compressed of its initial length L0